Find the sum of the first 12 terms of the arithmetic sequence whos general term is an=3n+ 5

Answer by dkppathak(439)

Find the sum of the first 12 terms of the arithmetic sequence whos general term is an=3n+ 5
a1=3x1+5=8
a2=3x2+5=11
a3=3x3+5=14
.
.
.
a12=3x12+5=41
therefor it will be AP whose 12 therms are as 8,11,14,--------41
a=8 d=3 n=12 l=41
sn=n(a+l)/2=12(8+41)/2=12x49/2=6x49=294
sum of 12 terms is 294